speed boat took 3 hours to travel the distance of 60 km Up aRiver against the current return trip took two hours find the average speed of the boat in Stillwater and the speed of the current
Answer by ankor@dixie-net.com(22740)

A speed boat took 3 hours to travel the distance of 60 km Up a River against the current.
Return trip took two hours.
Find the average speed of the boat in Still water and the speed of the current.
:
let s = boat speed in still water
let c = speed of the current
then
(s-c) = effective speed up the river
and
(s+c) = effective speed down river
:
Write a distance equation for each way; dist = time * speed
3(s-c) = 60
2(s+c) = 60
Simplify both equations, divide the 1st by 3, divide the 2nd by 2 and you have
s - c = 20
s + c = 30
------------adding eliminates c, find s
2s = 50
s = 50/s
s = 25 mph is the boat speed in still water
Find c, use the equation s + c = 30, replace s with 25
25 + c = 30
c = 30 - 25
c = 5 mph is the speed of the current
:
:
Check these solutions in the 1st original equation
3(25 - 5) = 60
3(20) = 60